### v2

# OpenMMReasoner: Pushing the Frontiers for Multimodal Reasoning with an Open and General Recipe

Abstract
Recent advancements in large reasoning models have fueled growing interest in extending such capabilities to multimodal domains. However, despite notable progress in visual reasoning, the lack of transparent and reproducible data curation and training strategies remains a major barrier to scalable research. In this work, we introduce OpenMMReasoner, a fully transparent two-stage recipe for multimodal reasoning spanning supervised fine-tuning (SFT) and reinforcement learning (RL). In the SFT stage, we construct an 874K-sample cold-start dataset with rigorous step-by-step validation, providing a strong foundation for reasoning capabilities. The subsequent RL stage leverages a 74K-sample dataset across diverse domains to further sharpen and stabilize these abilities, resulting in a more robust and efficient learning process. Extensive evaluations demonstrate that our training recipe not only surpasses strong baselines but also highlights the critical role of data quality and training design in shaping multimodal reasoning performance. Notably, our method achieves a 11.6% improvement over the Qwen2.5-VL-7B-Instruct baseline across nine multimodal reasoning benchmarks, establishing a solid empirical foundation for future large-scale multimodal reasoning research. We open-sourced all our codes, pipeline, and data at https://github.com/EvolvingLMMs-Lab/OpenMMReasoner.
Figure 1: Performance Comparison with State-of-the-Art Large Multimodal Reasoning Models across Various Benchmarks. Our proposed OpenMMReasoner consistently outperforms competing methods, highlighting its effectiveness in complex reasoning tasks.
11footnotetext: Equal contribution. Corresponding author.
1 Introduction
With the rapid progress of reinforcement learning with verifiable rewards (RLVR), leading models such as DeepSeek-R1 [10] and OpenAI o3 [25] have demonstrated strong reasoning capabilities across mathematics, programming, and scientific tasks. These advancements highlight RLVR as an effective paradigm for improving structured reasoning in large language models (LLMs). Motivated by this progress, researchers have increasingly explored RL-enhanced reasoning from LLMs to large multimodal models (LMMs). Recent studies [24, 4, 7, 37] push the boundaries of multimodal reasoning, showing that reinforcement learning (RL) can strengthen fine-grained visual understanding and complex cross-modal problem-solving. These developments suggest that the benefits of RLVR can transfer beyond text, offering a scalable path toward more capable and reliable multimodal reasoning systems.
Despite these advancements, transparency across the training pipeline remains limited. While numerous studies investigate the supervised-finetuning (SFT) and RL stages [3, 25, 6], few provide details of their data curation processes or conduct comprehensive ablation analyses. This lack of openness restricts reproducibility and obscures a deeper understanding of how reasoning-capable LMMs are actually built and how their training dynamics evolve. Recent efforts such as ThinkLite-VL [35] and MM-Eureka [24] propose clearer methodologies for RL data construction and offer useful insights for algorithm design. Other work, including Open Vision Reasoner (OVR) [37], attempts to address both SFT and RL, but does not provide a scalable, unified recipe that generalizes across tasks and modalities. Similarly, LLM-focused efforts [9] emphasize SFT while offering limited discussion of RL or multimodal extensions. Collectively, these limitations point to a critical gap: the lack of a transparent, scalable, and training recipe that unifies SFT and RL for building multimodal reasoning systems.
In this work, we present a comprehensive and scalable empirical study on training Large Multimodal Reasoning Models (LMRMs) from open-source LMMs (e.g., Qwen2.5-VL [3]). Through extensive experimentation, we develop simple yet effective data pipelines for constructing high-quality SFT and RL datasets, enabling reliable cold-start training of a strong LMRM. Our analysis provides detailed insights into which components most effectively contribute at each stage, revealing practical principles for building robust, generalizable, and efficient multimodal reasoning systems. Furthermore, we explore how to maximize scaling efficiency across both SFT and RL phases and summarize key lessons learned throughout the process.
Building on these insights, we introduce OpenMMReasoner, a fully transparent training recipe for multimodal reasoning. It features (1) a high-quality 874k-sample SFT dataset that establishes a strong reasoning foundation prior to RL, and (2) a 74k-sample RL dataset accompanied by detailed analysis of diverse designs and optimization strategies, which further sharpens and stabilizes these capabilities. As shown in Fig.˜1, our recipe yields a highly capable LMRM that consistently outperforms state-of-the-art methods such as OVR across a wide range of multimodal reasoning benchmarks, demonstrating both the effectiveness and scalability of our approach. The complete pipeline, spanning both the SFT and RL phases, is illustrated in Fig.˜2. As summarized in Tab.˜1, all components of our workflow are fully open and transparent, providing a comprehensive and reproducible view of the entire data curation and training process.
To summarize, our contributions are threefold: (1) Comprehensive Data Curation Insights. We present the first systematic study on curating high-quality SFT and RL data for multimodal reasoning, supported by extensive and rigorous experiments across diverse modalities and reasoning types. We found that scaling data diversity is a critical factor for curating high-quality datasets. While diversity in data sources is important, diversity in answers represents an additional essential axis for improvement. (2) A Strong SFT Recipe for Reasoning. We present a robust and reproducible SFT recipe that effectively equips LMMs with strong reasoning capabilities. Our approach incorporates step-by-step validation and offers practical insights for scaling a high-quality data pipeline focused on reasoning. By carefully selecting an appropriate teacher model for rejection sampling, and incorporating cross-domain data sources, we construct a cold-start dataset that exhibits both high diversity and quality, forming a solid reasoning foundation for subsequent RL. (3) An Advanced RL Recipe for Reasoning Enhancement. We conduct a comprehensive comparative analysis of multiple RL strategies, including GSPO [50], GRPO [30], and DAPO [44], to evaluate their stability, efficiency, and scaling behavior. By selecting the most suitable algorithm, we establish a robust RL pipeline that further sharpens and stabilizes reasoning abilities, delivering both high stability and superior performance.
Method Data Pipeline SFT Data RL Data Model
Qwen2.5-VL-7B-Instruct [3]
M2-Reasoning [2]
MiMo-VL [32]
OpenVisionReasoner [37]
[…]
Table 1: Comparison of Extent of Open-Sourcing across Existing LMRMs. OpenMMReasoner is the first study to fully open-source its data pipeline, SFT/RL datasets, model weights, enabling fully transparent reproduction.
Figure 2: Data Pipelines of OpenMMReasoner. We propose two training recipes covering both the SFT and RL phases. The pipeline begins by collecting diverse data sources and selecting teacher models to generate new answer traces. During the RL phase, we explore different algorithm choices and filtering strategies, leading to our final optimized recipe.
2 Related Work
RL has emerged as a powerful approach for enhancing reasoning in LLMs. Recent systems such as OpenAI’s o1 [26] and DeepSeek-R1 [10] have shown that multi-step reasoning and self-verification can emerge purely from large-scale RL. Building on these advances, recent works [24, 7, 37, 4, 17, 28, 40, 45, 42] extend RL-enhanced reasoning to multimodal models, while SFT studies [40, 41, 18] highlight the importance of high-quality supervision.
Despite this progress, many prior works [37, 1] lack transparent and reproducible training pipelines, and related efforts [9] focus only on textual reasoning. This limits reproducibility and obscures how data curation and training design affect outcomes. To address this, we present OpenMMReasoner, a scalable and fully open recipe covering both SFT and RL, offering practitioners clear guidance for building reliable multimodal reasoning models.
3 Supervised Fine-tuning Recipe
In this section, we present the data curation pipeline used to construct our SFT dataset. We begin with the initial data sources and describe how the dataset evolves through three stages: 1) data sourcing and formatting (103k raw questions), 2) data distillation and scaling (583k verified general-reasoning traces), and 3) cross-domain mixing to form the final 874k SFT recipe. The following experiments examine each stage and show how these design choices influence model performance. The distribution of the dataset across sources and domains is illustrated in Fig.˜3.
Figure 3: Data Source Distribution OpenMMReasoner. Our dataset comprises diverse sources across multiple domains, aiming to balance data diversity and efficiency for optimal performance.
SFT Datasets Benchmarks
Teacher Model Average MathVision MathVerse MathVista
[…]
Qwen2.5-VL-72B-Instruct 49.8 28.1 49.7 71.5
Qwen3-VL-235B-Instruct 50.5 29.3 52.9 69.4
Table 2: Teacher models improve performance and data efficiency. Even with limited data, leveraging a teacher model significantly enhances the model’s reasoning ability at scale.
SFT Datasets Benchmarks
Sampling Strategy Average MathVision MathVerse MathVista
$\times 1$ sampling 50.5 29.3 52.9 69.4
$\times 2$ sampling 52.7 30.8 54.4 72.9
$\times 4$ sampling 52.9 30.8 55.3 72.6
$\times 8$ sampling 55.2 34.6 57.1 73.7
Table 3: Repeated teacher sampling provides a scaling axis. With the same question sources, increasing answer diversity improves data quality, showing that both question diversity and answer diversity are important for model performance.
3.1 Data Curation
The overall pipeline is illustrated in Fig.˜2. The dataset is constructed through the following stages: 1) Data Sourcing, 2) Data Distillation and Scaling, and 3) Domain Mixing.
Data Sourcing.
We begin with approximately 103k raw question–answer pairs collected from public datasets including LLaVA-CoT [40], OpenVLThinker [7], and We-Math2.0 [27]. These samples mainly cover general VQA and reasoning-oriented tasks. Additional sources used later in the domain mixing stage (Section 3.5) include MMR1 [48] for image-based mathematical reasoning and MiroMind-M1 [21] for text-based mathematical reasoning.
Data Formatting.
Data sourced from different benchmarks often follow inconsistent answer styles and reasoning formats, which may introduce training instability [43, 21]. We standardize all samples to a unified reasoning format, normalizing textual structure and ensuring consistent step-wise outputs before entering the distillation stage.
3.2 Experimental Setup
For the SFT training, we adopt online packing and the Liger-Kernel [11] to accelerate training efficiency. All experiments are conducted using the LMMs-Engine [22] as the training framework and LMMs-Eval [19, 48] for standardized evaluation. We use Qwen2.5-VL-7B-Instruct [3] as our initial checkpoint to start, which also serves as our baseline. Each model is trained until convergence, and further implementation details are provided in the Appendix.
SFT Datasets Benchmarks
Filter Average MathVision MathVerse MathVista
No Filter 55.2 34.6 57.1 73.7
Length Filter 54.2 33.0 56.0 73.6
Difficulty Filter 51.3 30.6 53.3 70.2
Table 4: Over-filtering answers is detrimental. Although filtering is commonly used in SFT, it reduces dataset diversity and does not improve overall performance.
3.3 Data Distillation
Teacher Model Selection.
Prior work [10, 9] shows that distillation from a stronger model improves data quality and data efficiency. To identify the most suitable teacher, we compare answer traces distilled from multiple candidates, applying both format validation and answer verification. Only samples whose final answers pass a rule-based validator and an LLM-as-judge check are retained, producing around 59k verified traces. As shown in Tab.˜2, all teacher-based variants outperform the baseline with a clear leading, confirming that a stronger teacher yields higher-quality supervision. Both of the stronger teacher models provide an average gain of at least 4.5 points across all benchmarks, with Qwen3-VL-235B-Instruct [29] achieving the highest performance and selected as the teacher for distillation. This stage establishes the foundation for larger-scale answer sampling.
Scaling Data with verifiable answer.
[…]
SFT Datasets Benchmarks
Domain-mixing Average MathVision MathVerse MathVista
No Mix 55.2 34.6 57.1 73.7
ImgMath Mix 55.6 34.3 57.5 74.9
TxtMath Mix 55.6 35.3 57.2 74.2
Img+TxtMath 56.3 36.6 57.7 74.8
Table 5: Cross-domain data improves reasoning ability. Mixing data from multiple domains enhances the model’s reasoning performance, demonstrating effective reasoning transfer across domains.
3.4 Length and Difficulty Filtering
We next examine whether further data refinement improves performance. Two common filtering strategies—difficulty-based filtering and length-based filtering—are evaluated. Difficulty is approximated from previous sampling accuracy, while token count is used to remove extremely short examples. As shown in Tab.˜4, both filtering strategies reduce performance, which we attribute to the loss of answer diversity. Since filtering decreases sample variety without improving consistency, we adopt a no-filtering policy for the final recipe, preserving the full 583k general-reasoning set produced from the distillation stage.
Table 6: Evaluation Results on Visual Reasoning Benchmarks. Best results are bold and the second-best are underlined for open-source models. Indicates results reproduced by ourselves.
Model SFT Data RL Data MathVista MathVision MathVerse DynaMath WeMath LogicVista MMMU MMMU-Pro CharXiv
testmini test testmini testmini worst loose test val standard vision reas. desc.
Close-source Models
OpenAI-GPT-4o [14] - - 59.9 31.1 - 40.6 34.5 - 64.4 - - - - -
GPT-4o mini [14] - - 55.1 27.3 - 30.0 31.6 48.8 41.4 - 37.6 - 34.1 74.9
SFT Methods
LLaVA-OneVision-7B [20] 4.8M - 62.6 17.6 - 17.6 9.0 - 32.0 - 24.1 - 23.6 48.7
InternVL3-8B [51] - - 70.5 28.6 - 33.9 23.0 - 43.6 - - - 37.6 73.6
Qwen2.5-VL-7B [3] - - 69.2 25.5 25.6 41.1 21.8 53.1 47.9 51.8 37.9 35.1 36.4 67.3
LLaVA-OneVision-1.5-8B [2] 105M - 69.6 25.6 21.7 46.3 19.8 49.4 45.8 55.4 37.4 25.2 37.0 74.1
[…]
RL-based Methods
VLAA-Thinker-Qwen2.5-7B [4] 126k 25k 68.0 26.4 - 48.2 22.4 - 48.5 - - - - -
ThinkLite-7B-VL [35] - 11k 71.6 24.6 - 42.9 16.5 - 42.7 - - - - -
VL-Rethinker-7B [34] - 39k 73.7 28.4 - 46.4 17.8 - 42.7 - 41.7 - - -
M2-Reasoning [1] 6.2M 102k 75.0 42.1 - 40.4 - - 50.6 - - - - -
MMR1 [17] 1.6M 15k 72.0 31.8 29.0 55.4 27.9 68.0 48.9 52.4 41.1 37.1 43.5 71.1
OpenVLThinker-7B [7] 3.3k 9.6k 65.3 23.0 26.9 38.1 16.8 61.9 44.5 55.1 39.7 38.4 41.0 69.2
MM-Eureka-Qwen-7B [24] - 15.6k 72.6 28.1 32.1 45.4 23.0 59.8 46.3 54.4 40.1 37.1 42.4 74.1
OVR-7B [38] 2M 300k 72.1 51.8 38.2 54.6 33.5 64.8 54.8 51.8 50.2 29.1 44.5 73.6
[…]
3.5 Domain Mixing
To further enhance reasoning generalization, we incorporate supervised data from additional reasoning domains. While the 583k distilled dataset already provides strong multimodal reasoning, its coverage of mathematical reasoning remains limited. We therefore integrate MMR1 [17] (image-based math reasoning) and MiroMind-M1 [21] (text-based math reasoning). As shown in Tab.˜5, adding both types of mathematical supervision consistently improves performance across multimodal and reasoning benchmarks. Combining the general SFT data with these additional mathematical datasets yields our final 874k mixed SFT dataset.
3.6 Analysis and Insights
In summary, the dataset evolves through three major stages—from 103k raw questions, to 583k distilled general-reasoning samples, and finally to an 874k mixed SFT dataset incorporating both general and mathematical reasoning. As shown in Tab.˜6, our results across nine reasoning benchmarks [27, 27, 23, 39, 46, 47, 36, 52] show that our method achieves superior performance and data efficiency compared to other SFT approaches, demonstrating the effectiveness of our recipe in building a strong reasoning base model.
Our empirical findings highlight four key observations: (1) Answer diversity enhances reasoning. Increasing the diversity of generated answers consistently improves the model’s overall reasoning performance, even when using the same question sources, suggesting that exposure to varied solutions strengthens understanding. (2) Teacher model selection is crucial. Distilling from a strong teacher model substantially boosts the model’s reasoning ability while maintaining high data efficiency. Careful selection for teacher model directly affects the quality of the distilled dataset and the final model performance. (3) Over-filtering reduces diversity and performance. The best results are achieved without excessive filtering, indicating that maintaining greater answer diversity encourages more robust reasoning abilities. (4) Cross-domain knowledge improves generalization. Incorporating diverse data from multiple domains consistently enhances the model’s overall reasoning capabilities across tasks.
4 Reinforcement Learning Recipe
To further enhance the model’s generalization and strengthen its multimodal reasoning capabilities, we introduce an RL phase tailored for multimodal reasoning tasks. Building on the strong reasoning foundation established during the SFT stage, this phase serves to further sharpen and stabilize these abilities. In this section, we provide key insights into our algorithmic design and the strategies employed to ensure stable and efficient RL dynamics.
4.1 Preliminaries of Reinforcement Learning
Let each data pair $(q,a)$ be i.i.d from a distribution $\mathcal{D}$, where $q$ is a query and $a$ is the ground-truth answer. Given an LLM policy $\pi_{\theta}(\cdot|\cdot)$, let $o$ be an LLM-generated response to $q$, and $r(\cdot,\cdot)$ is a predefined reward function that quantifies whether the response $o$ yields $a$. RL-based fine-tuning aims to maximize this expected reward over $\mathcal{D}$, i.e.,
$\max_{\theta}\mathcal{J}(\pi_{\theta})\triangleq\mathbb{E}_{(q,a)\sim\mathcal{D}}\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}[r(o,a)].$ (1)
Group Relative Policy Optimization (GRPO)[30] is an efficient variant of PPO that removes the need for a critic network and Generalized Advantage Estimation (GAE), thereby reducing both memory usage and computational overhead. GRPO normalizes rewards within each rollout group to reduce variance and incorporates likelihood-ratio clipping with a KL-divergence penalty to constrain $\pi_{\theta}$ close to the initial SFT policy. The GRPO objective is defined as:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\begin{subarray}{c}\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid x)\end{subarray}}$ (2)
$\biggl[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\biggl(\min\Bigl(r_{i,t}(\theta)\hat{A}_{i,t}$
$\operatorname{clip}\left(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\Bigl)\biggr]$
Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) [44] addresses several limitations of GRPO, including entropy collapse, training instability, and length bias caused by sample-level loss. To mitigate these issues, DAPO introduces a decoupled clipping mechanism and a dynamic sampling strategy, leading to a more stable and balanced optimization objective:
$\mathcal{J}_{\text{DAPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\begin{subarray}{c}\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid x)\end{subarray}}$ (3)
$\biggl[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\biggl(\min\Bigl(r_{i,t}(\theta)\hat{A}_{i,t}$
$\operatorname{clip}\left(r_{i,t}(\theta),1-\varepsilon_{low},1+\varepsilon_{high}\right)\hat{A}_{i,t}\Bigl)\biggr]$
subject to $0<|\{o_{j}\mid\text{is\_equivalent}(a,o_{i})\}|<G$
Group Sequence Policy Optimization (GSPO) [50] tackles the token-level importance bias inherent in GRPO by introducing a sequence-level importance ratio for optimization. Additionally, GSPO employs a smaller clipping threshold $\varepsilon$ to enhance training stability. The resulting optimization objective for GSPO is defined as:
$\mathcal{J}_{\text{GSPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\begin{subarray}{c}\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid x)\end{subarray}}$ (4)
$\biggl[\frac{1}{G}\sum_{i=1}^{G}\biggl(\min\Bigl(s_{i}(\theta)\hat{A}_{i,t}$
$\operatorname{clip}\left(s_{i}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\Bigl)\biggr]$
where $s_{i}$ is the importance ratio based on sequence likelihood.
Table 7: Recipe selection results of different RL training strategies and coldstart starting point on Visual Reasoning Benchmarks. Unless otherwise specified, the rollout temperature is set to 1.0 by default.
Method Avg MMMU MMMU_Pro MathVista MathVerse MathVision CharXiv LogicVista WeMath DynaMath
val standard vision testmini test testmini test reas desc test loose worst
[…]
Baseline 49.4 54.4 39.3 37.3 74.8 57.7 33.9 36.6 39.7 76.1 46.2 67.2 29.3
DAPO + $\times 8$ rollout 45.6 53.3 37.7 34.6 71.7 51.0 28.6 27.9 40.7 73.0 44.4 58.3 26.4
DAPO + $\times 16$ rollout 48.9 52.9 40.6 35.6 74.3 56.2 30.3 34.5 43.4 77.0 46.9 67.6 27.4
GRPO + $\times 16$ rollout 51.1 54.6 42.8 39.4 77.1 58.3 33.9 37.1 43.1 73.8 51.1 70.2 32.1
GSPO + $\times 8$ rollout 51.6 54.9 41.2 38.8 76.9 61.4 35.9 39.9 45.0 73.8 46.9 71.7 32.3
GSPO + $\times 16$ rollout 54.3 57.8 44.1 40.6 79.5 63.8 38.8 43.6 46.1 73.5 50.0 79.0 34.9
GSPO + $\times 16$ rollout + curriculum 52.5 58.7 42.8 42.4 76.3 61.7 35.2 41.4 45.5 72.9 49.8 75.4 28.3
[…]
Baseline 49.2 55.6 40.6 37.7 73.7 57.1 35.2 34.6 39.5 74.1 48.9 69.1 24.0
DAPO + $\times 16$ rollout 49.6 54.8 42.0 37.6 74.1 57.8 36.8 33.9 39.1 74.3 48.7 71.6 24.0
DAPO + $\times 16$ rollout + temp. 1.4 49.3 56.9 41.0 37.3 73.3 57.8 33.2 32.2 40.1 75.0 51.6 70.2 23.4
GSPO + $\times 16$ rollout 51.0 55.1 42.4 39.6 75.1 59.4 36.2 35.4 44.6 73.0 49.1 74.0 27.9
GSPO + $\times 16$ rollout + temp. 1.4 7.4 13.4 3.9 28.3 3.4 22.0 0.0 0.0 0.2 0.0 8.3 0.2 9.2
[…]
Baseline 46.7 54.2 42.6 36.0 71.5 49.7 29.9 28.2 39.0 73.1 47.8 63.9 25.0
GRPO + $\times 8$ rollout 47.6 55.6 41.5 36.7 73.5 53.0 29.6 28.1 42.3 72.0 44.9 69.1 24.6
DAPO + $\times 8$ rollout 49.2 54.4 50.8 38.4 74.5 55.7 27.6 28.3 43.2 72.0 48.7 69.0 27.9
DAPO + $\times 8$ rollout + curriculum 47.0 56.0 40.5 36.8 71.6 51.2 32.9 28.1 36.2 73.6 46.2 66.8 24.6
4.2 Dataset Curation
Dataset Sourcing
Similar to the SFT stage, we begin by collecting diverse data samples across multiple domains. Our sources include MMEureka [24], ViRL [34], TQA [15], We-Math [28], PuzzleVQA [5], AlgoPuzzleVQA [8], and ThinkLiteVL [35], covering domains such as science, mathematics, charts, and puzzles.
Dataset Cleaning
To ensure answer validity, we extract and verify the final answers from each dataset. We then deduplicate by computing both image and text similarities to remove redundant questions. The cleaned datasets are merged to form the final dataset, resulting in approximately 74k samples for the RL stage. Overall data statistics are shown in Fig.˜3. Unless otherwise specified, all RL experiments are conducted using this dataset.
4.3 Experimental Setup
For the RL training stage, we utilize verl [31] and vllm [16] to accelerate the training process. The evaluation protocol remains consistent with that of the SFT stage, employing LMMs-Eval [19, 48] to ensure a unified and standardized evaluation setup. By default, we use temperature 1.0 and the top-performing checkpoint from the SFT phase as the starting point for RL, unless stated otherwise. The checkpoint corresponds to the best-performing model reported in Tab.˜5, where cross-domain mixing of image and text-based math data were applied, along with $\times 8$ sampling for answer traces.
Figure 4: Overall results across different algorithms. We conduct a systematic comparison of various algorithms under identical multimodal RL training settings. GSPO demonstrates the highest training stability, exploration capability, and overall efficiency.
Figure 5: Training dynamics on the validation set during RL. During RL training, we observe that textual reasoning ability improves alongside visual reasoning, even when trained solely on multimodal data, indicating strong cross-domain generalization of reasoning capabilities.
4.4 Training Recipe
RL Algorithm Selection
To identify the most suitable RL algorithm for multimodal reasoning, we compare GSPO [50], DAPO [44], and GRPO [30] under a unified setting. We evaluate their training dynamics in terms of stability, exploration, and efficiency, as illustrated in Fig.˜4. GSPO demonstrates faster convergence, higher rewards, and more stable behavior than DAPO and GRPO, achieving an effective balance between exploration and stability. While DAPO applies online filtering to remove zero-variance samples, it exhibits early entropy collapse and slower progress due to larger rollout requirements. GRPO shows moderate stability but converges more slowly. Based on these observations and the ablation results summarized in Sec.˜4.1, we adopt GSPO as the RL algorithm in our final training setup.
Reward Function
To balance task accuracy and output formatting, we adopt a composite reward function similar to those in [10, 13], which combines both aspects in a weighted manner. Specifically, the final reward $R$ for each sample is defined as:
$R=(1-\lambda_{\text{fmt}})\cdot R_{\text{acc}}+\lambda_{\text{fmt}}\cdot R_{\text{fmt}},$ (5)
where $R_{\text{acc}}$ measures the model’s correctness on the task objective, and $R_{\text{fmt}}$ captures the consistency of the answer format. The coefficient $\lambda_{\text{fmt}}\in[0,1]$ controls the trade-off between task accuracy and format adherence. We use $\lambda_{\text{fmt}}=0.1$ throughout our RL experiments.
Figure 6: Token efficiency comparison with OVR. OpenMMReasoner achieve better accuracy while using significantly less token budget.
Sampling Strategies Although RL algorithms naturally produce stronger learning signals on more challenging problems, their training efficiency remains limited. Previous work [33] employs a difficulty-based sampler, starting with easy tasks and gradually shifting to harder ones. We first use Qwen3-VL-8B [29] to track the pass rate of each problem as a difficulty metric and implement the following strategy: RL training begins with simpler tasks and progressively transitions to more complex ones. Ablation results in Sec.˜4.1 show that curriculum learning does not outperform the original mixed sampling approach; therefore, we adopt the no-sampling strategy in our final configuration.
Balancing Reasoning Capability and Efficiency.
Our empirical findings show that while OpenVisionReasoner [37] achieves competitive performance, its response length becomes excessively long, raising concerns about reasoning efficiency. This observation highlights an important question: how can a model maintain strong reasoning performance while remaining computationally efficient and adaptable across different problem types. In our reinforcement learning implementation, we adopt a similar overlength penalty strategy as proposed in DAPO [44] to mitigate the overthinking behavior and achieve a balanced trade-off between reasoning depth and efficiency. As shown in Fig.˜6, we evaluate models on two benchmarks, MMMU [46] and We-Math [27]. Our model demonstrates higher reasoning efficiency compared to OVR [37]. As shown in Fig.˜6, OVR requires excessively long reasoning trajectories to reach correct answers, whereas our model achieves a favorable balance between accuracy and efficiency, maintaining a reasonable reasoning budget while delivering superior overall performance.
4.5 Analysis and Insights
Overall Results.
We present the evaluation results of our final models for both the SFT and RL stages in Tab.˜6. Building on the strong reasoning foundation established during the SFT stage, the RL phase further sharpens and stabilizes these capabilities, leading to improved and more consistent performance. After RL, our model achieves state-of-the-art results on benchmarks such as WeMath [27], MathVerse [49], and MathVista [23] and demonstrating consistent improvement compare to SFT.
Textual Reasoning Transfers Alongside Strengthened Multimodal Reasoning.
As the model’s overall reasoning ability improves through RL training, we observe the gradual emergence of textual reasoning behaviors, suggesting a transfer of reasoning competence from multimodal to purely linguistic domains. As shown in Fig.˜5, validation performance on AIME24, AIME25, and AMC23 steadily increases throughout training, reflecting continuous and measurable gains in text-based reasoning capabilities. For AIME24 and AIME25, the reported accuracy represents the average over eight rollout runs. Prior work [12] has demonstrated that enhancements in mathematical reasoning can positively transfer to other reasoning domains. Evaluation results in Tab.˜8 compare baseline and RL-trained models, showing that textual reasoning improves and strengthens in tandem with multimodal reasoning across all training stages. These findings extend previous observations, indicating that cross-domain reasoning skills acquired via multimodal RL can effectively transfer to purely textual tasks, further highlighting the shared cognitive foundations and underlying mechanisms across different modalities.
Benchmarks
Model Average AIME24 AIME25 GPQA Diamond
Baseline 15.1 6.7 6.7 31.8
[…]
Table 8: Text reasoning ability transfer alongside visual reasoning improvement. As the model’s overall reasoning capability increases, its text-based reasoning ability also improves. Across different training stages, we observe significant gains in textual reasoning, indicating strong cross-domain generalization.
Key Factors for Stable Training Dynamics.
We identify two factors that critically affect RL training stability: rollout temperature and rollout count. First, higher rollout temperatures (e.g., 1.4) cause significant instability and occasional divergence, suggesting that excessive exploration amplifies policy gradient variance and destabilizes optimization. Second, the number of rollouts per update is central to maintaining convergence stability. To assess its impact, we compare $\times 8$ and $\times 16$ rollout configurations under both GRPO and DAPO. As shown in Fig.˜7, the $\times 16$ configuration consistently yields higher rewards and smoother dynamics. This effect is especially pronounced for DAPO, where the $\times 8$ setting exhibits severe late-stage instability: entropy initially fluctuates and then abruptly spikes, ultimately causing training collapse. Interestingly, although larger rollout counts seem more computationally expensive, actual wall-clock time remains comparable between $\times 8$ and $\times 16$ due to similar token-length constraints. Based on these observations, we adopt a moderate temperature and a $\times 16$ rollout configuration as our default setup.
Figure 7: Effect of rollout count on multimodal RL training stability. DAPO becomes unstable with fewer rollouts, while increasing the rollout count leads to more stable training dynamics.
In summary, our key observations from the RL phase are as follows: 1) GSPO outperforms other algorithms. GSPO demonstrates superior stability and faster convergence compared to alternative methods in multimodal RL training. 2) Token efficiency is crucial. While increasing reasoning steps at test time can improve performance, excessive tokens reduce efficiency. Our results show that a smaller reasoning budget can achieve comparable or even better accuracy. 3) Reasoning ability transfers across domains. Gains in reasoning during training consistently translate into stronger performance across multiple domains.
5 Conclusion
In this work, we present OpenMMReasoner, a transparent, scalable framework for training LMRMs via unified SFT and RL stages. Our study systematically examines how data curation, sampling strategies, and RL design shape multimodal reasoning. Experiments show that (1) well-designed SFT data, even in limited quantity, builds a strong reasoning foundation, and (2) carefully curated RL datasets, combined with effective algorithms like GSPO, enhance reasoning stability and performance. Scaling data diversity across domains and reasoning traces proves more valuable than merely increasing size. Structured sampling and difficulty-aware curricula improve efficiency, while well-defined reward signals strengthen reasoning precision and multimodal consistency. These insights highlight the importance of transparent, reproducible pipelines, and we hope OpenMMReasoner serves as a solid empirical foundation and open-source reference for scalable multimodal reasoning research.
Supplementary Material
Figure 1: Rollout Analysis over RL. With the progress of RL training, the model response contains more reflection word ratio.
Training Component SFT RL
optimizer AdamW AdamW
scheduler cosine constant
learning rate 5e-5 1e-6
weight decay 0.0 0.1
Training Steps 4300 1232
Warmup Steps 430 25
Max Length 61440 32792
Dynamic Bsz True True
Remove padding True True
Liger Kernel True False
Table 1: Detail parameters for SFT and RL training.
1 Implementation Details
Table 2: Additional evaluation results on Sampling Strategy. Scaling up with more answer generation leads to better result.
Method Avg MMMU MMMU_Pro MathVista MathVerse MathVision CharXiv LogicVista WeMath DynaMath
val standard vision testmini test testmini test reas desc test loose worst
[…]
$\times 1$ sampling 46.7 54.2 42.6 36.0 71.5 49.7 29.9 28.2 39.0 73.1 47.8 63.9 25.0
$\times 2$ sampling 47.5 54.4 39.1 34.6 72.9 54.4 30.9 30.8 37.5 73.6 49.6 69.3 23.0
$\times 4$ sampling 48.2 56.9 41.0 37.6 72.6 55.3 30.3 30.8 40.7 75.4 45.8 69.3 23.0
$\times 8$ sampling 49.2 55.6 40.6 37.7 73.7 57.1 35.2 34.6 39.5 74.1 48.9 69.1 24.0
Table 3: Reward ablation result.
Method Avg MMMU MMMU_Pro MathVista MathVerse MathVision CharXiv LogicVista WeMath DynaMath
val standard vision testmini test testmini test reas desc test loose worst
[…]
$\lambda_{fmt}=0.1$ 51.1 54.6 42.8 39.4 77.1 58.3 33.9 37.1 43.1 73.8 51.1 70.2 32.1
$\lambda_{fmt}=0.3$ 47.0 51.3 36.7 33.9 73.9 55.8 29.6 34.1 36.5 75.9 43.3 64.3 29.1
$\lambda_{fmt}=0.5$ 45.0 48.3 34.3 32.4 74.2 53.1 26.0 27.9 39.8 75.6 41.1 58.1 29.3
$\lambda_{fmt}=0.7$ 48.8 53.4 37.8 35.8 74.8 56.8 32.9 35.8 39.4 75.7 46.9 66.9 29.1
System Prompt for model You are a helpful assistant. When the user asks a question, your response must include two parts: first, the reasoning process enclosed in <think>…</think> tags, then the final answer enclosed in <answer>…</answer> tags. Please provide a clear, concise response within <answer> </answer> tags that directly addresses the question.
Table 4: The prompt that used in evaluation.
System Prompt for judge model You are a strict evaluator assessing answer correctness. You must output 1 for fully correct answers and 0 for any other case. You will receive the question, the ground truth answer, and the model prediction. # Input Question: “‘ {question} “‘ Ground Truth Answer: “‘ {answer} “‘ Model Prediction: “‘ {prediction} “‘ # Evaluation Rules - For multiple-choice questions: Score 1 if the predicted answer matches the ground truth answer, it can be directly in option letters or the content of the options. - For open-ended questions: * Score 1 if the prediction matches the answer semantically, it can be in different format. * Score 0 for partially correct answers or answers with extra incorrect information, even if the reasoning process is correct. - Ignore minor differences in formatting, capitalization, or spacing since the model may explain in a different way. - Treat numerical answers as correct if they match within reasonable precision - For questions requiring units, both value and unit must be correct # Strict Output format 1 or 0
Table 5: The prompt that used in evaluation.
1.1 Training Details
We describe the training procedures for our best-performing models in both the SFT and RL stages.
SFT.
To maximize training throughput and reduce memory consumption, we apply online stream packing with an iterable dataset, removing padding tokens to avoid unnecessary computation. We use a packing length of 61,440 tokens. For each batch, we dynamically compute the input token lengths of incoming samples and fill the batch buffer until it is fully packed. Because packing is performed online, the total number of epochs cannot be predetermined; instead, we train each model until convergence. The full set of experimental hyperparameters is provided in  Tab.˜1.
RL.
For RL training, we use a global batch size of 128, which strikes a balance between on-policy stability and training speed. Log probabilities are computed using a dynamic batch-size implementation without padding to further reduce memory usage. During generation, we set the maximum number of new tokens to 28,696 and cap the prompt length at 4,096 tokens, resulting a max length at 32792. We use a temperature of 1.0 and keep this configuration fixed across all experiments. Due to the high computational cost of RL training, we run training until the reward saturates. The detailed hyperparameters are listed in  Tab.˜1.
1.2 Evaluation Details
We now describe our evaluation setup for multimodal reasoning benchmarks. The SFT and RL models share the same evaluation configuration. We use the system prompt shown in  Tab.˜4 to ensure the model outputs both the reasoning trace and the final answer in an extractable format. The extracted answer is then validated using a two-stage process: a rule-based validator followed by an LLM-as-judge validator. We first apply the rule-based validator to minimize evaluation cost; if the answer cannot be verified, we fall back to the LLM-as-judge, using the prompt provided in  Tab.˜5. For all evaluations, we use a temperature of 0.0 for reproducibility and set the maximum generation length to 49K tokens, except for AIME, where we use a temperature of 1.0. We employ vLLM [16] as the serving engine to accelerate inference.
2 Additional Result and Analysis
Sampling Scaling Results.
We present the full evaluation results for different sampling strategies in  Sec.˜1. As shown in the table, scaling up the sampling strategy improves the average score from 46.7 to 49.2, demonstrating the effectiveness of increasing answer diversity along this axis.
Rollout Analysis.
During RL training, we record all rollout logs and analyze the proportion of reflection-related words in the model’s outputs. We observe that as the reward increases over training steps, the proportion of reflection words also rises. To illustrate this, we generate a word cloud from the final-step rollout outputs—after removing noise words—and find that reflection cues such as “let,” “wait,” and “think” appear frequently in the model’s responses. This indicates that our RL recipe effectively encourages the model to engage in more explicit reasoning as its capabilities improve. The results are shown in  Fig.˜1.
Reward $\lambda_{fmt}$ Ablation.
To assess the impact of the $\lambda_{fmt}$ parameter in RL, we conduct additional experiments using four values: 0.1, 0.3, 0.5, and 0.7, under the same experimental setup as the GRPO configuration. The evaluation results are presented in  Sec.˜1. We observe that a lower format-reward weight, specifically $\lambda_{fmt}=0.1$, consistently yields the best performance. Consequently, we adopt $\lambda_{fmt}=0.1$ in our final configuration.
3 Examples
Data examples.
We present several examples in Tab.˜6 and Tab.˜7 from our reasoning dataset to demonstrate the high quality of the answer traces generated by our method. These cases illustrate how the model decomposes complex questions into structured reasoning steps and produces coherent, verifiable conclusions.
Qualitative Results.
We present several qualitative examples in Tab.˜8 and Tab.˜9 to demonstrate the effectiveness and robustness of our model. These examples highlight the model’s ability to accurately interpret complex multimodal inputs, generate coherent reasoning steps, and produce reliable answers across diverse scenarios, showcasing both its generalization capability and practical utility.
4 Limitation and Future Work
Our work primarily focuses on a single model family—Qwen2.5-VL-Instruct—and evaluates performance mainly within the image domain. While our approach demonstrates strong gains in multimodal reasoning, it does not extend to other modalities such as video or audio, limiting its applicability to a broader set of real-world tasks. Additionally, although we explore scaling strategies in both SFT and RL stages, we have not yet identified the upper bound of model performance under further scaling, leaving open the question of how far the current recipe can be pushed.
For future work, we aim to extend our methodology to a wider range of domains, including video, audio, and richer temporal-stream modalities. Another promising direction is to explore generation capabilities across multiple modalities simultaneously, enabling more coherent and context-aware multimodal reasoning and content synthesis. Broadening the model family used for evaluation and training will further help validate the generality and robustness of our findings across architectures and modality configurations.
ColdStart Data Samples
Question As shown in the figure, line EF passes through the intersection point O of the diagonals of the parallelogram and intersects AD and BC at points E and F, respectively. If the area of the parallelogram is 12, the sum of the areas of triangles AOE and BOF equals ( )A. 2 B. 3 C. 4 D. Cannot be determined Output the thinking process in <think> </think> and final answer in <answer> </answer> tags.
Answer <think>In a parallelogram, the diagonals bisect each other, so point O is the midpoint of both diagonals AC and BD. Also, since EF passes through O and intersects AD and BC at E and F respectively, we can use symmetry and properties of parallelograms. Because O is the center of the parallelogram, any line through O divides the parallelogram into two regions of equal area. Moreover, triangles AOE and COF are congruent (by ASA or SAS, since AO = OC, ∠AOE = ∠COF, and ∠OAE = ∠OCF due to parallel lines and transversal), and similarly, triangles BOF and DOE are congruent.Therefore, the area of triangle AOE equals the area of triangle COF, and the area of triangle BOF equals the area of triangle DOE.The total area of the parallelogram is 12. The diagonals divide it into 4 triangles of equal area: AOB, BOC, COD, DOA — each of area 3.Now, consider the two shaded triangles: AOE and BOF. Since EF passes through O, and by symmetry, the sum of the areas of AOE and BOF is equal to the area of triangle AOB (which is one of the four equal-area triangles formed by the diagonals).Why? Because:- Area(AOE) + Area(BOF) = Area(AOB)This is because the parts not included in AOE and BOF (namely, EOD and FOC) are congruent to AOE and BOF respectively, so swapping them doesn’t change the total.Alternatively, since the entire parallelogram is divided into 4 triangles of area 3 each, and AOE + BOF together make up exactly one of those triangles (AOB), their combined area is 3.Thus, the answer is 3.</think><answer>B. 3</answer>
Table 6: An example of the coldstart data sample.
ColdStart Data Samples
Question What is the man doing with the beans?A. cookingB. eating themC. countingD. selling themAnswer with the option’s letter from the given choices directly. Output the thinking process in <think> </think> and final answer in <answer> </answer> tags.
Answer <think>The image shows a man sitting beside a cart loaded with green beans in what appears to be a busy street market. He is handling the beans and seems to be interacting with potential customers, which is typical behavior for a vendor. There is no visible cooking equipment, nor is he eating or counting the beans. The context strongly suggests he is selling them.</think><answer>D</answer>
Table 7: An example of the coldstart data sample.
Qualitative Data Sample
Question A. Political instability leading to population decline B. The spread of pathogens across the Silk Road C. Development of new trade routes D. Climate change affecting the Silk Road E. Migrations to areas of Central Asia for resettlement F. Technological advancements in transportationG. Invasions by Mongol tribes H. Large-scale famine due to crop failures I. Economic prosperity and population growth J. Rise of religious conflicts along the Silk Road Answer the question in the image.
Answer B
Ours <think>The graph shows population trends for Roman and Han populations over time, with a general decline in population for both over centuries. This decline is most plausibly explained by widespread disease, as the Silk Road facilitated the spread of pathogens across regions, leading to population reductions. While political instability or other factors could contribute, the most direct and historically supported cause for such a widespread population decline across empires is the spread of pathogens.</think><answer>B. The spread of pathogens across the Silk Road</answer>
Qwen2.5-VL-7B-Instruct <think>The graph shows a general decline in population for both the Roman and Han empires over the period from 200 C.E. to 600 C.E. This trend is not consistent with economic prosperity, technological advancements, or large-scale famine, which would typically lead to population growth. The spread of pathogens across the Silk Road could potentially cause a decline in population, but the graph does not provide specific information about the cause of the decline. The rise of religious conflicts along the Silk Road could also lead to population decline, but the graph does not provide specific information about the cause of the decline. The development of new trade routes and the rise of religious conflicts along the Silk Road are not directly related to the population decline shown in the graph. The graph does not provide specific information about the cause of the decline, but it is reasonable to infer that the decline in population could be due to political instability, which could lead to a decline in the overall population.</think><answer>A. Political instability leading to population decline</answer>
Table 8: An example of our model inference result
Qualitative Data Sample
Question A. <image 3> B. <image 2> C. <image 4> D. <image 5> E. <image 1> Answer the question in the image.
Answer D
Ours <think> The relationship between arousal and performance on a challenging task is typically described by the Yerkes-Dodson Law, which states that performance increases with arousal up to a point, after which further increases in arousal lead to a decline in performance. This forms a bell-shaped curve 2014 peak performance at moderate arousal, with performance decreasing as arousal rises too high or falls too low. Looking at the options: - (A) Inverted V: performance improves with low arousal, then drops 2014 doesn2019t match. - (B) Straight line decreasing: performance always decreases with arousal 2014 too linear and wrong direction. - (C) Straight line increasing: performance always increases with arousal 2014 incorrect. - (D) Bell curve: performance peaks at moderate arousal 2014 matches Yerkes-Dodson Law. - (E) U-shape: performance improves with low arousal, then drops 2014 opposite of what2019s expected. Thus, (D) best represents the relationship. The question asks to choose from A-E, and (D) is labeled as image 5 in the options mapping. So, the correct answer is D. </think><answer>D. <image 5></answer>
Qwen2.5-VL-7B-Instruct <think>The relationship between a person’s level of arousal and their performance on a challenging task is often described by the Yerkes-Dodson Law, which suggests that performance increases with arousal up to a certain point and then decreases. This relationship is typically represented by a bell-shaped curve, indicating an optimal level of arousal for peak performance.</think><answer>E</answer>
Table 9: An example of our model inference result